OpenStudy (dancergirl45):

Which statement is true about the sum square root of 50 + square root of 18?

OpenStudy (dancergirl45):

It is rational and equal to 2 rational and equal to 8 irrational and equal to 2 to the square route of 2 irrational and equal to 8 to the square route of 2

OpenStudy (anonymous):

it's definitely irrational

OpenStudy (dancergirl45):

ok so C or D

OpenStudy (math&ing001):

\[\sqrt{50}+\sqrt{18}=\sqrt{2*25}+\sqrt{2*9}=\sqrt{2}*\sqrt{25}+\sqrt{2}*\sqrt{9}=5\sqrt{2}+3\sqrt{2}=8\sqrt{2}\]

OpenStudy (dancergirl45):

so D?

OpenStudy (math&ing001):

Yeah

OpenStudy (dancergirl45):

Thank you guys so much!!

OpenStudy (dancergirl45):

I will open a question can u help me with a few more